# On Entropy Control in LLM-RL Algorithms

### v1

@@ 1 Introduction
RL seeks to optimize the reward received by a sequential decision making system. In recent years, RL has proven to be an effective tool for training LLMs (Yang et al., 2025; DeepSeek-AI, 2025; Comanici et al., 2025). The advances of LLMs in math, coding and planning tasks has been astonishing, with their performance on competitive benchmarks drastically increasing after RL training.
@@
However, it is observed that entropy regularization offers little gains in LLM-RL training. Specifically, the experimental results to be shown in Section 5 suggest that entropy-regularized GRPO yields minimal gain compared to basic GRPO. In addition, Cui et al. (2025) observes that the validation accuracy is unchanged under different scaling of the entropy bonus in LLM-math tasks. These results are particularly underwhelming compared to those in other deep RL tasks including robotics and games, where the benefit of entropy bonus is significant (see, e.g., (Haarnoja et al., 2018, Figure 3)). Moreover, such empirical contradiction also indicates a theoretical gap between the existing analysis which justifies the entropy’s benefit (Mei et al., 2020) and its effect in LLM-RL. Therefore, a careful study and a remedy for this issue is in dire needs, as the potential gain from entropy bonus is yet to be unlocked for LLM training.
@@
  • A theory on the entropy effect and its issues in LLM-RL. Under no entropy bonus, we show that entropy collapse indicates learning stagnancy and give a performance bound. Then we show that entropy regularization can fail to improve this result due to its self-induced bias which increases with the sparsity of optimal responses in LLM tasks.
@@ 1.1 Related works
Entropy regularization in RL. Entropy-regularized RL was initially introduced in (Williams & Peng, 1991). It has been commonly used in various popular policy-based deep RL algorithms (Mnih et al., 2016; Schulman et al., 2017; Haarnoja et al., 2018) which have provided ample empirical evidence for its effectiveness in robotic and games tasks. Entropy regularization’s optimization benefits have also been empirically (Ahmed et al., 2019) and theoretically (Mei et al., 2020) studied. However, it does not give notable performance gains for LLMs (see, e.g., Section 5 and (Cui et al., 2025; Klein et al., 2023)). As a result, alternative entropy control techniques are often adopted. In (Zhang et al., 2024) reshapes the reward function to regulate the policy. Or in a concurrent work (Cui et al., 2025), the algorithm clips or regulates the parts the policy update that decrease entropy too much. To our best knowledge, existing works do not answer the question of why and when entropy regularization can fail in LLM-RL, and have not uncovered the potential benefits of entropy bonus.
@@ 2 Preliminaries
Finite-horizon Markov decision process (MDP). In LLM-RL setting, the learning task can be modeled as a finite-horizon MDP defined by a $\mathcal{M}=\{{\mathcal{S}},{\mathcal{A}},{\mathcal{P}},r,H\}$, where $\mathcal{S}$} is a finite state space (e.g., input space of the LLM), $\mathcal{A}$} is a finite action space (e.g., LLM’s vocabulary), ${\mathcal{P}}(s_{t+1}=\cdot|s_{t},a_{t})$ is the state transition kernel, which is assumed to be deterministic in LLM tasks (e.g., $s_{t+1}$ is a concatenation of $(s_{t},a_{t})$). Function $r(s,a)\in[0,1]$ assigns a reward to the pair $(s,a)$. Horizon $H$ corresponds to the max response length. An LLM-policy parameterized by $\theta\in{\mathbb{R}}^{d}$ is denoted as $\pi_{\theta}(a|s)$, which assigns a probability for each token $a\in{\mathcal{A}}$ given input $s\in{\mathcal{S}}$.
@@
where $\pi_{\theta}(s_{t})\coloneqq\pi_{\theta}(\cdot|s_{t})$, the expectation is taken over the trajectory $(a_{t},s_{t+1},\dots,a_{H-1})$ where $a_{t}\sim\pi_{\theta}(s_{t})$ for each $t$. Similarly, we can define the Q-function as $Q_{h}^{\pi_{\theta}}(s,a)=r(s,a)+V_{h+1}^{\pi_{\theta}}(s^{\prime})$ with $s^{\prime}=(s,a)$, and the advantage function as $A_{h}^{\pi_{\theta}}(s,a)=Q_{h}^{\pi_{\theta}}(s,a)-V_{h}^{\pi_{\theta}}(s)$. Given a dataset $\mathcal{D}$} containing input queries, the objective of RL is to maximize the expected value on $\mathcal{D}$}:
@@
Entropy-regularized RL. Given $\mathcal{D}$}, we can define the entropy of the policy $\pi_{\theta}$ as
@@
In maximum entropy RL, we optimize for the entropy regularized objective:
$\max_{\theta}V_{\lambda}^{\pi_{\theta}}({\mathcal{D}})\coloneqq V^{\pi_{\theta}}({\mathcal{D}})+\lambda{\mathcal{H}}(\pi_{\theta})=\mathbb{E}_{\pi_{\theta}}\Big{[}\sum_{t=0}^{H-1}r(s_{t},a_{t})-\lambda\log\pi_{\theta}(a_{t}|s_{t})\Big{]}.$ (2.4)
We can then analogously define the entropy-regularized value functions given the initial state and action $s,a$ and initial time-step $h$ as $V_{h,\lambda}^{\pi_{\theta}}(s)\coloneqq\mathbb{E}_{\pi_{\theta}}\big{[}\sum_{t=h}^{H-1}\big{(}r(s_{t},a_{t})-\lambda\log\pi_{\theta}(a_{t}|s_{t})\big{)}|s_{h}=s\big{]}$ and $Q_{h,\lambda}^{\pi_{\theta}}(s,a)\coloneqq r(s,a)+V_{h+1,\lambda}^{\pi_{\theta}}(s^{\prime})\text{ with }s^{\prime}=(s,a)$. Then the entropy-regularized advantage function is defined as $A_{h,\lambda}^{\pi_{\theta}}(s,a)=Q_{h,\lambda}^{\pi_{\theta}}(s,a)-\lambda\log\pi_{\theta}(a|s)-V_{h,\lambda}^{\pi_{\theta}}(s)$.
2.1 Policy-gradient based LLM-RL algorithms
We give a brief review over some prominent policy optimization algorithms. The gradients of the objectives are essentially estimates of $\nabla V^{\pi_{\theta}}({\mathcal{D}})$, thus making them policy-gradient based methods.
PPO-clip family. Given the sampling policy $\pi_{\rm b}$, the objectives of PPO-clip algorithms (Schulman et al., 2017) can be written as
${\mathcal{L}}_{\rm clip}(\theta)\!=\!\mathbb{E}_{s_{0}\sim{\mathcal{D}},\{a_{t}\sim\pi_{\rm b}(s_{t})\}_{t\leq H-1}}\!\Big{[}\min\Big{(}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\rm b}(a_{t}|s_{t})}\hat{A}_{t},{\rm clip}\Big{(}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\rm b}(a_{t}|s_{t})},1\!-\!\epsilon_{\rm low},1\!+\!\epsilon_{\rm high}\!\Big{)}\!\hat{A}_{t}\Big{)}\Big{]}$ (2.5)
where $\hat{A}_{t}$ is an estimate of the advantage $A_{t}^{\pi_{\theta}}(s_{t},a_{t})$. GRPO uses a trajectory-level Monte-Carlo estimate of the advantage. DAPO additionally decouples the clip ratio by setting different $\epsilon_{\rm low},\epsilon_{\rm high}$ and incorporates extra sampling constraints and overlong response penalty.
REINFORCE with a baseline. With $\pi_{b}=\pi_{\theta}$, REINFORCE objective can be written as
${\mathcal{L}}_{\rm reinforce}(\theta)\!=\!\mathbb{E}_{s_{0}\sim{\mathcal{D}},\{a_{t}\sim\pi_{b}(s_{t})\}_{t\leq H-1}}\!\Big{[}\hat{A}_{t}\log\pi_{\theta}(a_{t}|s_{t})\Big{]}$ (2.6)
where the gradient calculation w.r.t. $\theta$ ignores $\pi_{b}$. Some recent works (Ahmadian et al., 2024; Chu et al., 2025) found simple REINFORCE can outperform PPO-type algorithms in certain settings.
@@ 3 A theory on entropy effect in policy gradient based LLM-RL
where $\theta_{s,a}$ is the logit of token $a$ given input $s$. The LLM-RL algorithms without entropy regularization are generally guaranteed to converge to an $\epsilon$-stationary point of the RL objective $V^{\pi_{\theta}}({\mathcal{D}})$ satisfying $\|\nabla V^{\pi_{\theta}}({\mathcal{D}})\|\leq\epsilon$ (Agarwal et al., 2021; Jin et al., 2023). When doing policy optimization without regularization, (Cui et al., 2025) observes that the policy entropy quickly diminishes as performance increases, and ultimately performance saturates when entropy completely collapses. In the following result, we give some theoretical insights into this case.
@@
Assume the policy is a softmax. Then the following two statements hold.
@@
    $V^{\pi^{*}}(s_{0})-V^{\pi_{\theta}}(s_{0})\leq|{\mathcal{D}}|\frac{1}{C^{\pi_{\theta}}(s_{0})}\epsilon$
    where $\pi^{*}\in\arg\max_{\pi}V^{\pi}({\mathcal{D}})$, $C^{\pi_{\theta}}(s_{0})\coloneqq H^{-0.5}\max_{(a_{0},\dots,a_{H-1})\in{\mathcal{A}}_{H}^{*}(s_{0})}\Pi_{t=0}^{H-1}\pi_{\theta}(a_{t}|s_{t})$ in which ${\mathcal{A}}_{H}^{*}(s_{0})=\{(a_{0},a_{1},\dots,a_{H-1})\in{\mathcal{A}}^{H}~|~\exists\pi^{*},\Pi_{t=0}^{H-1}\pi^{*}(a_{t}|s_{t})>0\}$ is the set of all optimal responses given query $s_{0}$.
Note $\big{\|}\nabla V^{\pi_{\theta}}({\mathcal{D}})\big{\|}$ is the policy gradient without entropy regularization.
The first bullet (I) suggests the policy entropy is an indicator of the policy stationarity, that is, a small entropy indicates a small policy gradient $\|\nabla V^{\pi_{\theta}}({\mathcal{D}})\big{\|}$ and the convergence of the policy. The second bullet (II) quantifies the actual performance of the almost stationary policy, where the reward on query $s_{0}$ is bounded by $\mathcal{O}(\epsilon/C^{\pi_{\theta}}(s_{0}))$. The factor $C^{\pi_{\theta}}(s_{0})$ can be controlled (bounded away from 0) when the initial LLM and the RL algorithm can sufficiently explore the optimal response to $s_{0}$. For example, one can either use a large batch size (Klein et al., 2023) or a strong initial model (Weissmann et al., 2024) to control $C^{\pi_{\theta}}(s_{0})$. In this case, the performance is ultimately bounded by $\mathcal{O}(\epsilon)$. The error $\epsilon$ decreases with prolonged RL training, while it usually cannot decrease to 0 due to the presence of sampling noise or the advantage estimate error.
@@
$V^{\pi^{*}}(s_{0})-V^{\pi_{\theta}}(s_{0})\leq|{\mathcal{D}}|^{2}\frac{1}{C_{\lambda}^{\pi_{\theta}}(s_{0})}\frac{\epsilon^{2}}{2\lambda}+\underbrace{\lambda H\log\frac{|{\mathcal{A}}|}{|{\mathcal{A}}_{H}^{*}(s_{0})|^{\frac{1}{H}}}}_{\rm bias}$
@@
Similar conditions to Propositions 1.(II)&2 have been derived in (Mei et al., 2020) for the discounted infinite horizon MDPs, while our results hold for the finite horizon MDPs under a deterministic transition. Proposition 2 also provides a more accurate bound for the entropy bias.
Entropy regularization suffers from large bias due to immense response space with sparse optimality in LLM tasks. As compared to no entropy control case in Proposition (II), the above bound’s dependence on $\epsilon$ improves to $\mathcal{O}(\epsilon^{2}/2\lambda)$. However, this optimization benefit does not come free as a bias term is introduced. The entropy bias is $\mathcal{O}(H\log(|{\mathcal{A}}|/|{\mathcal{A}}_{H}^{*}(s_{0})|^{\frac{1}{H}}))$, which increases with the response space size $H\log|{\mathcal{A}}|$ and the sparsity of optimal responses $\log(1/|{\mathcal{A}}_{H}^{*}(s_{0})|)$. The bias is especially ubiquitous in LLM-RL, where the response space is typically extremely large (hundreds of thousands tokens to choose from in each step) as compared to that in, e.g., classic control and games where the action space size and horizon are typically at the hundreds (Brockman et al., 2016; Silver et al., 2017).
To leverage the benefit of entropy regularization, it is crucial to reduce the entropy bias. In the following sections, we propose our recipe for this issue and empirically demonstrate its effectiveness.
@@ 4 AEnt: Adaptive entropy regularization with token space clamping
In this section, we will first introduce the two core components of our method, and then present the AEnt algorithm.
@@ 4.1 Entropy with token space clamping
Recall that maximum entropy RL maximizes $V^{\pi_{\theta}}({\mathcal{D}})+\lambda{\mathcal{H}}(\pi_{\theta})$ where
@@
Entropy ${\mathcal{H}}(\pi)$ is maximized by the uniform policy $\pi_{\rm uniform}(a|s)=1/|{\mathcal{A}}|$. Maximizing the entropy pulls the LLM policy towards $\pi_{\rm uniform}(a|s)=1/|{\mathcal{A}}|$, increasing the likelihood of low-probability actions while decreasing those of the high-probability ones. Intuitively, this helps when the optimal actions have low probabilities and are thus less likely to be sampled and reinforced. Such mechanism works well in the RL tasks where the discrete action space is small (Brockman et al., 2016). While it is extremely inefficient in LLM-RL setting since $\mathcal{A}$} is prohibitively immense with sparse optimal tokens. Specifically, when $1/|{\mathcal{A}}|$ is small, pulling $\pi_{\theta}(a|s)$ for every $a\in{\mathcal{A}}$ towards $1/|{\mathcal{A}}|$ gives weak gains and produces large bias due to the large amount of non-optimal tokens in the complete token space.
Figure 1: Policy after 1-step policy optimization (PO) with no entropy control, the traditional entropy bonus and the clamped entropy bonus, respectively. The numbers result from direct computation.
@@
The clamped entropy is evaluated by a re-normalized policy $\tilde{\pi}_{\theta}$ on a size-reduced, input-dependent token space ${\mathcal{A}}(s)$. By the theoretical insights in Section 3, regularizing on a smaller response space with denser optimality generally leads to reduced bias. With this principle, we set ${\mathcal{A}}(s)$ as the the top-probability tokens set of $\pi_{\theta}(s)$. The intuition is that since the base models are pre-trained or fine-tuned prior to the RL phase, the bottom probability tokens are unlikely to be optimal. We find leaving them out reduces entropy-induced bias and leads to notable performance gains in LLM-RL.
To directly observe the effect of clamped entropy, we provide a numerical illustration in Figure 1. We can see that policy optimization without entropy control tends to reinforce the high-probability locally optimal tokens, while squeezing down the low-probability optimal ones that are less likely to be sampled. This creates a negative-effect loop that cannot be reversed once policy entropy collapses. On the other hand, using the clamped entropy bonus lifts up the low-probability optimal tokens. This makes them more likely to be sampled in subsequent iterations and get properly reinforced.

### v2

@@ 1 Introduction
RL seeks to maximize the reward received by a sequential decision making system. In recent years, RL has proven to be an effective tool for training LLMs (Yang et al., 2025; DeepSeek-AI, 2025; Comanici et al., 2025). The advances of LLMs in math, coding and planning tasks has been astonishing, with their performance on competitive benchmarks drastically increasing after RL training.
@@
However, it is observed that entropy regularization offers little gains in LLM-RL training. Specifically, the experimental results to be shown in Section 5 suggest that entropy-regularized GRPO yields minimal gain compared to basic GRPO. In addition, Cui et al. (2025) observes that the validation accuracy is unchanged under different scaling of the entropy bonus in some LLM-math tasks. These results are particularly underwhelming compared to those in other deep RL tasks including robotics and games, where the benefit of entropy bonus is significant (see, e.g., (Haarnoja et al., 2018, Figure 3)). Moreover, such empirical contradiction also indicates a theoretical gap between the existing analysis which justifies the entropy’s benefit (Mei et al., 2020) and its effect in LLM-RL. Therefore, a careful study and a remedy for this issue is in dire needs, as the potential gain from entropy bonus is yet to be unlocked for LLM training.
@@
  • A theory on the entropy effect and its issues in LLM-RL. Under no entropy bonus, we show that entropy collapse indicates learning stagnancy and give a performance bound. Then we show that entropy regularization can fail to improve this result under LLM’s large response space and the task’s sparse optimality.
@@ 1.1 Related works
Entropy regularization in RL. Entropy-regularized RL was initially introduced in (Williams & Peng, 1991). It has been commonly used in various popular policy-based deep RL algorithms (Mnih et al., 2016; Schulman et al., 2017; Haarnoja et al., 2018) which have provided ample empirical evidence for its effectiveness in robotic and games tasks. Entropy regularization’s optimization benefits have also been empirically (Ahmed et al., 2019) and theoretically (Mei et al., 2020; Klein et al., 2023) studied. However, it does not give notable performance gains for LLMs (see, e.g., Section 5 and (Cui et al., 2025)). As a result, alternative entropy control techniques are often adopted. In (Zhang et al., 2024) reshapes the reward function to regulate the policy. Or in a concurrent work (Cui et al., 2025), the algorithm clips or regulates the parts the policy update that decrease entropy too much. To our best knowledge, existing works do not answer the question of why and when entropy regularization can fail in LLM-RL, and have not uncovered the potential benefits of entropy bonus.
@@ 2 Preliminaries
Finite-horizon Markov decision process (MDP). In LLM-RL setting, the learning task can be modeled as a finite-horizon MDP defined by a $\mathcal{M}=\{{\mathcal{S}},{\mathcal{A}},{\mathcal{P}},r,H\}$, where $\mathcal{S}$} is a finite state space (e.g., input token sequence of the LLM), $\mathcal{A}$} is a finite action space (e.g., LLM’s vocabulary), and the state transits by $s_{t+1}={\mathcal{P}}(s_{t},a_{t})$ where $\mathcal{P}$} is a concatenation operation of the input sequence $s_{t}$ and the output token $a_{t}$. Function $r(s,a)\in[0,1]$ assigns a reward to $(s,a)$. Horizon $H$ is the max response length. An LLM-policy parameterized by $\theta\in{\mathbb{R}}^{d}$ is denoted as $\pi_{\theta}(a|s)$, which assigns a probability for each token $a\in{\mathcal{A}}$ given input $s\in{\mathcal{S}}$.
@@
where $\pi_{\theta}(s_{t})\coloneqq\pi_{\theta}(\cdot|s_{t})$, the expectation is taken over the trajectory $(a_{t},s_{t+1},\dots,a_{H-1})$ where $a_{t}\sim\pi_{\theta}(s_{t})$ for each $t$. Given a dataset $\mathcal{D}$} containing input queries, the objective of RL is
@@
Entropy regularized RL. Given $\mathcal{D}$}, we can define the entropy of the policy $\pi_{\theta}$ as
@@
Entropy regularized RL aims to maximize the entropy regularized objective $V_{\lambda}^{\pi_{\theta}}({\mathcal{D}})\coloneqq V^{\pi_{\theta}}({\mathcal{D}})+\lambda{\mathcal{H}}(\pi_{\theta})$. Due to space limitation, we defer some definitions to Appendix A.1.
PPO-clip family. To solve for 2.2, a prominent algorithm is the PPO-clip (Schulman et al., 2017). Given the sampling policy $\pi_{\rm b}$, the objectives of PPO-clip algorithms can be written as
${\mathcal{L}}_{\rm PO}(\theta)\!=\!\mathbb{E}_{s_{0}\sim{\mathcal{D}},\{a_{t}\sim\pi_{\rm b}(s_{t})\}_{t\leq H-1}}\!\Big[\min\Big(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\rm b}(a_{t}|s_{t})}\hat{A}_{t},{\rm clip}\Big(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\rm b}(a_{t}|s_{t})},1\!-\!\epsilon_{\rm low},1\!+\!\epsilon_{\rm high}\!\Big)\!\hat{A}_{t}\Big)\Big]$ (2.4)
where $\hat{A}_{t}$ is an estimate of the advantage function. GRPO uses a trajectory-level Monte-Carlo estimate of the advantage. DAPO additionally decouples the clip ratio by setting different $\epsilon_{\rm low},\epsilon_{\rm high}$ and incorporates extra sampling constraints and overlong response penalty. Given a suitable clip range, the PPO-clip algorithm can be viewed as a policy gradient algorithm (Jin et al., 2023).
@@ 3 A theory on entropy effect in policy gradient based LLM-RL
where $\theta_{s,a}$ is the logit of token $a$ given input $s$. The policy gradient based algorithms without entropy regularization, including PPO-clip, are generally guaranteed to converge to an $\epsilon$-stationary point of the RL objective $V^{\pi_{\theta}}({\mathcal{D}})$ satisfying $\|\nabla V^{\pi_{\theta}}({\mathcal{D}})\|\leq\epsilon$ (Agarwal et al., 2021; Jin et al., 2023). When doing policy optimization without regularization, Cui et al. (2025) observes that the policy entropy quickly diminishes as performance increases, and ultimately performance saturates when entropy completely collapses. In the following result, we give some theoretical insights into this case.
@@
Assume the policy is a softmax. We have:
@@
    $V^{\pi^{*}}(s_{0})-V^{\pi_{\theta}}(s_{0})\leq\frac{1}{C^{\pi_{\theta}}(s_{0})}\epsilon$
    where $\pi^{*}\in\arg\max_{\pi}V^{\pi}({\mathcal{D}})$, $C^{\pi_{\theta}}(s_{0})\coloneqq\frac{1}{\sqrt{H}|{\mathcal{D}}|}\max_{(a_{0},\dots,a_{H-1})\in{\mathcal{A}}_{H}^{*}(s_{0})}\Pi_{t=0}^{H-1}\pi_{\theta}(a_{t}|s_{t})$ in which ${\mathcal{A}}_{H}^{*}(s_{0})=\{(a_{0},a_{1},\dots,a_{H-1})\in{\mathcal{A}}^{H}~|~\exists\pi^{*},\Pi_{t=0}^{H-1}\pi^{*}(a_{t}|s_{t})>0\}$ is the set of all optimal responses given query $s_{0}$.
The first bullet (I) suggests the policy entropy is an indicator of the policy stationarity, that is, a small entropy indicates a small policy gradient $\|\nabla V^{\pi_{\theta}}({\mathcal{D}})\big\|$ and the convergence of the policy. The second bullet (II) quantifies the actual performance of the almost stationary policy, where the reward on query $s_{0}$ is bounded by $\mathcal{O}(\epsilon/C^{\pi_{\theta}}(s_{0}))$. The factor $C^{\pi_{\theta}}(s_{0})$ can be controlled (bounded away from 0) when the initial LLM and the RL algorithm can sufficiently explore the optimal response to $s_{0}$. For example, one can either use a large batch size (Klein et al., 2023) or a strong initial model (Weissmann et al., 2024) to control $C^{\pi_{\theta}}(s_{0})$. In this case, the performance is ultimately bounded by $\mathcal{O}(\epsilon)$. The error $\epsilon$ decreases with prolonged RL training, while it usually cannot decrease to 0 due to the presence of sampling noise or the advantage estimation error.
Figure 1: Test in a controlled MDP with a large action space of size $|{\mathcal{A}}|=10^{5}$ and increasingly sparse optimal actions.
@@
$V^{\pi^{*}}(s_{0})-V^{\pi_{\theta}}(s_{0})\leq\frac{1}{C_{\lambda}^{\pi_{\theta}}(s_{0})}\frac{\epsilon^{2}}{2\lambda}+\lambda H\log\frac{|{\mathcal{A}}|}{|{\mathcal{A}}_{H}^{*}(s_{0})|^{\frac{1}{H}}}$
@@
Similar conditions to Propositions 1.(II)&2 have been derived in (Mei et al., 2020) for the discounted infinite horizon MDPs, while our results hold for the finite horizon MDPs under a concatenation transition. Proposition 2 also provides a more accurate bound for the entropy bias.
Entropy regularization suffers from immense response space with sparse optimality in LLM tasks. As compared to no entropy control case in Proposition (II), the above bound’s dependence on $\epsilon$ improves to $\mathcal{O}(\epsilon^{2}/2\lambda)$ when $C_{\lambda}^{\pi_{\theta}}(s_{0})$ is bounded away from 0. However, this optimization benefit does not come free as a bias term is introduced. The entropy bias is $\mathcal{O}(H\log(|{\mathcal{A}}|/|{\mathcal{A}}_{H}^{*}(s_{0})|^{\frac{1}{H}}))$, which increases with the response space size $H\log|{\mathcal{A}}|$ and the sparsity of optimal responses $\log(1/|{\mathcal{A}}_{H}^{*}(s_{0})|)$. The bias is especially ubiquitous in LLM-RL, where the response space is typically extremely large (hundreds of thousands tokens to sample from in each step) as compared to that in, e.g., classic control and games where the action space size and horizon are typically at the hundreds (Brockman et al., 2016; Silver et al., 2017). To better demonstrate this effect, we report a numerical test result in Figure 1. The test is done in a synthetic MDP with $|{\mathcal{A}}|=10^{5}$, and the optimal action is sparse with various numbers in $\{1,5,10,15\}$ (see Appendix A.5 for details). It is observed that entropy regularization leads to gains over no-regularization when number of optimal actions is $10,15$, but offers no gains when the optimal action becomes too sparse with fewer than $5$ optimal actions.
To unlock the benefit of entropy regularization, it is crucial to mitigate the negative effect caused by the large response space in LLM tasks. In the following sections, we propose our recipe for this issue and empirically demonstrate its effectiveness.
@@ 4 AEnt: Adaptive entropy regularization with token space clamping
In this section, we will first introduce the core components of our method and then the full algorithm.
@@ 4.1 Entropy with token space clamping
Recall the entropy regularized RL objective is $V^{\pi_{\theta}}({\mathcal{D}})+\lambda{\mathcal{H}}(\pi_{\theta})$ where
@@
Entropy is maximized by the uniform policy $\pi_{\rm uniform}(a|s)=1/|{\mathcal{A}}|$. Maximizing the entropy pulls the LLM policy towards $\pi_{\rm uniform}(a|s)=1/|{\mathcal{A}}|$, increasing the likelihood of low-probability actions while decreasing those of the high-probability ones. Intuitively, this helps when the optimal actions have low probabilities and are thus less likely to be sampled and reinforced. Such mechanism works well in the RL tasks where the discrete action space is small (Brockman et al., 2016). While it is extremely inefficient in LLM-RL setting since $\mathcal{A}$} is prohibitively immense with sparse optimal tokens. Specifically, when $1/|{\mathcal{A}}|$ is small, pulling $\pi_{\theta}(a|s)$ for every $a\in{\mathcal{A}}$ towards $1/|{\mathcal{A}}|$ gives weak gains and produces large bias due to the large amount of non-optimal tokens in the complete token space.
@@
The clamped entropy is evaluated by a re-normalized policy $\tilde{\pi}_{\theta}$ on a size-reduced, input-dependent token space ${\mathcal{A}}(s)$. By the insights from Proposition 2 and Figure 1, regularizing on a smaller response space with denser optimality generally leads to reduced bias and larger gains. With this principle, we set ${\mathcal{A}}(s)$ as the the top-probability tokens set of $\pi_{\theta}(s)$. The intuition is that since the base models are pre-trained or fine-tuned prior to the RL phase, the bottom probability tokens are unlikely to be optimal. We find leaving them out reduces entropy-induced bias and leads to performance gains. It can be observed from the toy demonstration in Figure 1 that clamped entropy regularization leads to performance gains when entropy regularization does not (number of optimal actions $\leq 5$), and is generally more robust to optimality sparsity increase.